# Generalization Error of the Tilted Empirical Risk

### v1

@@ D.1 Uniform bounds: details for unbounded loss
$\begin{split}&\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp{\gamma\ell(h,Z_{i})}\Big{)}\\ &\geq\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{1}{\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\left(\sqrt{\frac{4\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\frac{4\log(2/\delta)}{3n}\right)\\ &\geq\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{1}{(1-\zeta)\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\sqrt{\frac{4\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}\\ &\qquad-\frac{1}{(1-\zeta)\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\frac{4\log(2/\delta)}{3n}\\ &\geq\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{\exp(-\gamma R_{T}^{u})}{(1-\zeta)}\sqrt{\frac{4\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}-\frac{\exp(-\gamma R_{T}^{u})}{(1-\zeta)}\frac{4\log(2/\delta)}{3n}\\ &\geq\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{|\gamma|2\exp(-\gamma R_{T}^{u})}{(1-\zeta)}\sqrt{\frac{M_{2}^{u}\log(2/\delta)}{n}}-\frac{\exp(-\gamma R_{T}^{u})}{(1-\zeta)}\frac{4\log(2/\delta)}{3n}.\end{split}$ (71)
Here we used that by Assumption 4.1,
$\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\leq\exp[\mathbb{E}(\gamma\ell(h,\tilde{Z}))]\leq\exp(\gamma R_{T}^{u})$
@@
$\mathrm{Var}(\exp(\gamma\ell(h,\tilde{Z})))\leq[\mathbb{E}{\exp(\gamma\ell(h,\tilde{Z}))^{2}}]\leq\exp[\mathbb{E}{(\gamma\ell(h,\tilde{Z}))^{2}}]\leq\exp(\gamma^{2}M_{2}^{u}).$
@@
Now applying $\exp(\gamma R_{T}^{u})\leq\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\leq 1$ and $\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\leq\gamma^{2}M_{2}^{u}$ completes the proof. ∎
@@
Under the same assumptions in Theorem 4.6 and a finite hypothesis space, then for $n\geq\frac{(4\gamma^{2}M_{2}^{u}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma R_{T}^{u})}$, for $\gamma<0$ and with probability at least $(1-\delta)$, the absolute value of the titled generalization error satisfies $\begin{split}\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|&\leq\frac{2\exp(-\gamma R_{T}^{u})}{(1-\zeta)}\sqrt{\frac{M_{2}^{u}(\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma R_{T}^{u})(\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta))}{3n\gamma(1-\zeta)}-\frac{\gamma}{2}M_{2}^{u}.\end{split}$ (72)
@@
$\begin{split}|\mathrm{gen}_{\gamma}(h,S)|&\leq\frac{2\exp(-\gamma R_{T}^{u})}{(1-\zeta)}\sqrt{\frac{M_{2}^{u}(\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma R_{T}^{u})(\log(2/\delta))}{3n\gamma(1-\zeta)}-\frac{\gamma}{2}M_{2}^{u}.\end{split}$ (73)
Then, using the uniform bound Lemma B.8 completes the proof. ∎
D.2 Information-theoretic bounds: details for unbounded loss
Proposition 4.10.
Given Assumption 4.4 and Assumption 4.3, the following inequality holds for $\gamma<0$,
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\leq\begin{cases}\exp(-\gamma R_{T})\sqrt{\frac{M_{2}I(H;S)}{n}},\qquad&\frac{I(H;S)}{n}\leq\frac{\gamma^{2}M_{2}}{2}\\ \frac{\exp(-\gamma R_{T})}{|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}\Big{)},\qquad&\frac{I(H;S)}{n}>\frac{\gamma^{2}M_{2}}{2}.\end{cases}\end{split}$ (74)
@@ D.2 Information-theoretic bounds: details for unbounded loss
For $\gamma<0$ and we use that $0\leq\exp(\gamma\ell(H,Z_{i}))\leq 1$ and $\mathrm{Var}(\exp(\gamma\ell(H,Z_{i})))\leq\gamma^{2}\mathrm{Var}(\ell(H,Z_{i}))\leq\gamma^{2}M_{2}$. We also have $\mathrm{Var}(\exp(\gamma\ell(H,Z_{i})))\leq\frac{1}{4}$. Note that the variable $\exp(\gamma\ell(H,Z_{i}))$ is sub-exponential with parameters $(\gamma^{2}M_{2},1)$ under the distribution $P_{H}\otimes\mu$. Using the approach in (Bu et al., 2020; Aminian et al., 2021a) for the sub-exponential case, we have
$\begin{split}&\Big{|}\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{|}\\ &\leq\begin{cases}|\gamma|\sqrt{M_{2}\frac{I(H;S)}{n}}&\quad\mbox{ if }\frac{I(H;S)}{n}\leq\frac{\gamma^{2}M_{2}}{2}\\ \frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}M_{2}}{2}.\end{cases}\end{split}$ (75)
Therefore, we have for $\frac{I(H;S)}{n}\leq\frac{\gamma^{2}M_{2}}{2}$,
$\begin{split}\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\leq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+|\gamma|\sqrt{\frac{M_{2}I(H;S)}{n}}\Big{)}.\end{split}$ (76)
@@
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{|\gamma|}{\gamma\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\sqrt{\frac{M_{2}I(H;S)}{n}}.\end{split}$ (77)
For $\frac{I(H;S)}{n}>\frac{\gamma^{2}M_{2}}{2}$, we have
$\begin{split}&\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\\ &\leq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+\frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}\Big{)}.\end{split}$ (78)
@@
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{1}{\gamma\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}\Big{)}.\end{split}$ (79)
As under Assumption 4.3, we have $\exp(\gamma R_{T})\leq\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]$, the final result follows. ∎
Corollary 4.11.
Under Assumption 4.4 and Assumption 4.3, the following inequality holds for $\gamma<0$;
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\geq\begin{cases}-\exp(-\gamma R_{T})\sqrt{\frac{M_{2}I(H;S)}{n}},\qquad&\max\big{(}\frac{2I(H;S)}{\gamma^{2}M_{2}},\frac{\gamma^{2}M_{2}I(H;S)}{\exp(2\gamma R_{T})}\big{)}\leq n\\ \frac{\exp(-\gamma R_{T})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}\Big{)},\qquad&\min(\frac{\exp(\gamma R_{T})-\frac{\gamma^{2}M_{2}}{2}}{I(H;S)},\frac{2I(H;S)}{\gamma^{2}M_{2}})>n,.\end{cases}\end{split}$ (80)
@@
Similar to Proposition 4.10, we have
@@
$\begin{split}&\Big{|}\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{|}\\ &\leq\begin{cases}|\gamma|\sqrt{M_{2}\frac{I(H;S)}{n}}&\quad\mbox{ if }\frac{I(H;S)}{n}\leq\frac{\gamma^{2}M_{2}}{2}\\ \frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}M_{2}}{2}.\end{cases}\end{split}$ (81)
Therefore, we have for $\frac{I(H;S)}{n}\leq\frac{\gamma^{2}M_{2}}{2}$,
@@
$\begin{split}\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\geq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+\gamma\sqrt{\frac{M_{2}I(H;S)}{n}}\Big{)}.\end{split}$ (82)
Using $\log(y+x)\leq\log(y)+\frac{x}{y}$, we have
@@
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\leq\frac{\gamma}{\gamma\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\sqrt{\frac{M_{2}I(H;S)}{n}},\end{split}$ (83)
where it holds for $n\geq\frac{\gamma^{2}M_{2}I(H;S)}{\exp(2\gamma R_{T})}$. For $\frac{I(H;S)}{n}>\frac{\gamma^{2}M_{2}}{2}$, we have
@@
$\begin{split}&\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\\ &\geq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}-\frac{I(H;S)}{n}-\frac{\gamma^{2}M_{2}}{2}\Big{)}.\end{split}$ (84)
Using $\log(y+x)\leq\log(y)+\frac{x}{y}$, we obtain for $\frac{\exp(\gamma R_{T})-\frac{\gamma^{2}M_{2}}{2}}{I(H;S)}\geq n$,
@@
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{1}{\gamma\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}\Big{)}.\end{split}$ (85)
@@
As under Assumption 4.3, we have $\exp(\gamma R_{T})\leq\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]$, the final result follows. ∎
@@
Using Proposition 4.10 and Lemma B.3, we can derive the following upper bound on the expected generalization error.
Theorem 4.12.

### v2

@@ E.1 Uniform bounds: details for unbounded loss
$\begin{split}&\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp{\gamma\ell(h,Z_{i})}\Big{)}\\ &\geq\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{1}{\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\left(\sqrt{\frac{4\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\frac{4\log(2/\delta)}{3n}\right)\\ &\geq\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{1}{(1-\zeta)\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\sqrt{\frac{4\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}\\ &\qquad-\frac{1}{(1-\zeta)\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\frac{4\log(2/\delta)}{3n}\\ &\geq\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{4\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}-\frac{\exp(-\gamma\kappa_{u})}{(1-\zeta)}\frac{4\log(2/\delta)}{3n}\\ &\geq\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{|\gamma|2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{\log(2/\delta)}{n}}-\frac{\exp(-\gamma\kappa_{u})}{(1-\zeta)}\frac{4\log(2/\delta)}{3n}.\end{split}$ (72)
Here we used that by Assumption 4.1 and Lemma C.8,
$\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\leq\exp[\mathbb{E}(\gamma\ell(h,\tilde{Z}))]\leq\exp(\gamma\kappa_{u})$
@@
$\mathrm{Var}(\exp(\gamma\ell(h,\tilde{Z})))\leq[\mathbb{E}{\exp(\gamma\ell(h,\tilde{Z}))^{2}}]\leq\exp[\mathbb{E}{(\gamma\ell(h,\tilde{Z}))^{2}}]\leq\exp(\gamma^{2}\kappa_{u}^{2}).$
@@
Now applying $\exp(\gamma\kappa_{u})\leq\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\leq 1$ and $\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\leq\gamma^{2}\kappa_{u}^{2}$ completes the proof. ∎
@@
Under the same assumptions in Theorem 4.3 and a finite hypothesis space, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$, for $\gamma<0$ and with probability at least $(1-\delta)$, the absolute value of the titled generalization error satisfies $\begin{split}\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|&\leq\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{(\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma\kappa_{u})(\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta))}{3n\gamma(1-\zeta)}-\frac{\gamma}{2}\kappa_{u}^{2}.\end{split}$ (73)
@@
$\begin{split}|\mathrm{gen}_{\gamma}(h,S)|&\leq\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{(\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma\kappa_{u})(\log(2/\delta))}{3n\gamma(1-\zeta)}-\frac{\gamma}{2}\kappa_{u}^{2}.\end{split}$ (74)
Then, using the uniform bound Lemma C.11 completes the proof. ∎
E.2 Information-theoretic Bounds
In the information-theoretic approach for the unbounded loss function, we relax the uniform assumption, Assumption 4.1, as follows,
Assumption E.1 (Bounded second moment).
The learning algorithm $P_{H|S}$, loss function $\ell$, and $\mu$ are such that there is a constant $\kappa_{t}\in\mathbb{R}^{+}$ where the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\mathbb{E}_{P_{H}\otimes\mu}[\ell^{2}(H,Z)]\leq\kappa_{t}^{2}$.
Proposition E.2.
Given Assumption E.1, the following inequality holds for $\gamma<0$,
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\leq\begin{cases}\exp(-\gamma\kappa_{t})\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}},\qquad&\mbox{ if }\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{-\exp(-\gamma\kappa_{t})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)},\qquad&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}\end{cases}\end{split}$
@@ E.2 Information-theoretic Bounds
For $\gamma<0$ and we use that $0\leq\exp(\gamma\ell(H,\tilde{Z}))\leq 1$ and $\mathrm{Var}(\exp(\gamma\ell(H,\tilde{Z})))\leq\gamma^{2}\mathrm{Var}(\ell(H,\tilde{Z}))\leq\gamma^{2}\kappa_{t}^{2}$. Note that the variable $\exp(\gamma\ell(H,\tilde{Z}))$ is sub-exponential with parameters $(\gamma^{2}\kappa_{t}^{2},1)$ under the distribution $P_{H}\otimes\mu$. Using the approach in (Bu et al., 2020; Aminian et al., 2021a) for the sub-exponential case, we have
$\begin{split}&\Big{|}\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{|}\\ &\leq\begin{cases}|\gamma|\sqrt{\kappa_{t}^{2}\frac{I(H;S)}{n}}&\quad\mbox{ if}\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}.\end{cases}\end{split}$ (75)
Therefore, we have for $\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}$,
$\begin{split}\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\leq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+|\gamma|\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\Big{)}.\end{split}$ (76)
@@
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{|\gamma|}{\gamma\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}.\end{split}$ (77)
For $\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}$, we have
$\begin{split}&\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\\ &\leq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}.\end{split}$ (78)
@@
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{1}{\gamma\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}.\end{split}$ (79)
As under Assumption E.1 and Lemma C.8, we have $\exp(\gamma\kappa_{t})\leq\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]$, the final result follows. ∎
Corollary E.3.
Given Assumption E.1, the following inequality holds for $\gamma<0$ and some $\zeta\in(0,1)$,
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\geq\begin{cases}\frac{-\exp(-\gamma\kappa_{t})}{(1-\zeta)}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}},\quad&\mbox{ if }\max\big{(}\frac{2I(H;S)}{\gamma^{2}\kappa_{t}^{2}},\frac{\gamma^{2}\kappa_{t}^{2}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}\big{)}\leq n\\ \frac{\exp(-\gamma\kappa_{t})}{(1-\zeta)\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)},\quad&\mbox{ if }\min(\frac{4I(H;S)\big{(}4\zeta^{-2}-1\big{)}}{\gamma^{2}\kappa_{t}^{2}},\frac{2I(H;S)}{\gamma^{2}\kappa_{t}^{2}})>n.\end{cases}\end{split}$
@@
Similar to Proposition E.2, we have
@@
$\begin{split}&\Big{|}\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{|}\\ &\leq\begin{cases}|\gamma|\sqrt{\kappa_{t}^{2}\frac{I(H;S)}{n}}&\quad\mbox{ if}\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}.\end{cases}\end{split}$ (80)
Therefore, we have for $\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}$,
@@
$\begin{split}\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\geq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+\gamma\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\Big{)}.\end{split}$ (81)
Assume for now that there is a $\zeta\in(0,1)$ such that
@@
$|\gamma|\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\leq\zeta\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}$ (82)
As $\log(y-x)=\log(y)+\log(1-\frac{x}{y})\geq\log(y)-\frac{x}{y-x}$ for $y>x>0$, then by taking $y=\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}$ and $x=|\gamma|\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}$, so that with (71) we have $y-x\geq(1-\zeta)y>0,$ taking logarithms on both sides of (81) gives that
@@
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{1}{(1-\zeta)\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}},\end{split}$ (83)
where it holds for $n\geq\frac{\gamma^{2}\kappa_{t}^{2}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}$. For $\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}$, we have
@@
$\begin{split}&\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\geq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}-\frac{I(H;S)}{n}-\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}.\end{split}$ (84)
As $\log(y-x)=\log(y)+\log(1-\frac{x}{y})\geq\log(y)-\frac{x}{y-x}$, we obtain,
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{1}{\gamma(1-\zeta)\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)},\end{split}$ (85)
where it holds for $\frac{4I(H;S)\big{(}4\zeta^{-2}-1\big{)}}{\gamma^{2}\kappa_{t}^{2}}\geq n$. Note that, we have,
$\begin{split}&\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}\\ &\geq\frac{|\gamma|}{\zeta}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\\ &\geq\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}.\end{split}$ (86)
Therefore, we have,
$\frac{|\gamma|}{\zeta}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\geq\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2},$ (87)
where 87 is a quadratic inequality in $\sqrt{n}$ and using that $(a+b)^{2}\leq 2a^{2}+2b^{2}$ yields,
$\frac{4I(H;S)\big{(}4\zeta^{-2}-1\big{)}}{\gamma^{2}\kappa_{t}^{2}}\geq n.$
@@
As under Assumption E.1 and Lemma C.8, we have $\exp(\gamma\kappa_{t})\leq\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]$, the final result follows. ∎
@@
Using Proposition E.2, we can obtain the following upper bound on the expected generalization error.
Theorem E.4.